# Searching for GeV Gamma-Ray Polarization and Axion-Like Particles with AMS-02

### v1

@@ Abstract
We study the detectability of GeV-band gamma-ray polarization with the AMS-02 experiment and a proposed successor, from Galactic and extragalactic sources. Characterizing gamma-ray polarization in this energy range could shed light on gamma-ray emission mechanisms in the sources; physics beyond the Standard Model, such as the presence of axion-like particles (ALPs), could also induce a distinctive energy-dependent polarization signal due to propagation effects in magnetic fields. We present estimates for the minimum detectable polarization from bright sources and the forecast reach for axion-like particles (ALPs). We show that for an example bright extragalactic source (NGC1275), AMS-02 will only have sensitivity to ALP-induced polarization from currently-open parameter space if the B-field configuration is unusually favorable to a signal. However, a successor experiment such as AMS-100 would be expected to probe new parameter space even for pessimistic B-field models, with prospects to measure the energy-dependence of such a signal. For Galactic sources, polarization measurements could provide a unique test of scenarios where ALPs induce energy-dependent features in the photon intensity. However, in the absence of a bright transient source (such as a Galactic supernova), the parameter space that would be probed by this approach with ten years of AMS-100 data is already nominally excluded by other experiments, although this conflict may be avoided in specific ALP models.
@@ I Introduction
The existence of dark matter has long been one of the greatest puzzles in our understanding of the Universe. One of the leading candidates for dark matter particles is a hypothetical particle called the axion [1, 2] introduced by Peccei and Quinn [3, 4] to resolve the puzzle of the lack of observed CP violation in quantum chromodynamics (QCD). More generally, axion-like particles (ALPs) are ultralight pseudoscalar bosons predicted in many theoretical frameworks beyond the Standard Model of particle physics, e.g., [5, 6, 7]. In contrast to QCD axions, the characteristic parameters of ALPs—the mass $m_{a}$ and the photon-axion coupling $g_{a\gamma\gamma}$—are not necessarily related to each other. In the presence of an external magnetic field, the photon-ALP mixing can lead to energy-dependent modifications to both the flux and the polarization of photons from astrophysical sources, over a wide range of frequencies.
The effects of ALP dark matter on the polarization of light have been studied for various astrophysical sources (e.g., gamma-ray bursts [8], magnetic white dwarfs [9], active galactic nuclei [10, 11], and pulsars [12]). These studies are typically applied to photon energies below the MeV scale. Extending the studies of polarization to the GeV region requires the measurement of electron-positron pair production. To date, no data on polarization has been released for photons of energies above the pair creation threshold ($\sim$ 1 MeV), although the Fermi-LAT gamma-ray telescope has the potential to detect high degrees of linear polarization from some of the brightest gamma-ray sources [13]. Challenges and future directions for gamma-ray polarimetry have been reviewed in Ref. [14].
In the gamma-ray band, ALP-induced modification of intensity spectra from Galactic pulsars [15] and supernova remnants [16] has already been studied in data from the Fermi Large Area Telescope (Fermi-LAT). Intriguingly, these studies have claimed a hint of a possible signal, with a favored mass of $\sim$ neV and coupling $g_{a\gamma\gamma}\sim 10^{-10}$ GeV-1. The tension of these results with the limits from the CAST helioscope [17] might be reconciled by considering the different environmental conditions inside the Sun and the dilute interstellar medium where the photon-axion mixing occurs [18], or by non-standard ALP couplings [19]. Additional experimental data such as polarization from these Galactic sources could provide valuable input in testing the ALP hypothesis and perhaps probing the properties of ALPs.
@@
In this paper, we estimate the possible imprint of ALP dark matter in energy-dependent gamma-ray polarization, and discuss the implications for the sensitivity of AMS-02 and proposed next generation spectrometer experiments in space. Proposals for such experiments include ALADInO and AMS-100  [26, 27], with acceptance exceeding AMS-02 by roughly a factor of 10 or 100 respectively; we will show results for AMS-100 as a benchmark. We consider the extragalactic source NGC1275 at the center of the Perseus cluster, which has been studied in the context of ALP searches using X-ray polarimetry (e.g. in Ref. [11]) and gamma-ray intensity measurements (e.g. [28, 18, 29]), and a range of bright Galactic pulsars and supernovae.
@@ II.1 Photon-ALP mixing
$(E+\mathcal{M}-i\partial_{z})\begin{pmatrix}\Ket{\gamma_{x}}\\ \Ket{\gamma_{y}}\\ \Ket{a}\end{pmatrix}=0,$ (2)
where |$\Ket{\gamma_{x}}$ and |$\Ket{\gamma_{y}}$ correspond to the two linear polarization states of the photon field. We can always choose a coordinate system such that the external magnetic field transverse to the propagation direction, $B_{\perp}$, aligns with the $y$-direction. The mixing matrix is then [31, 32]
@@
$\begin{split}\Delta_{\text{pl}}&=-1.1\times 10^{-7}\left(\frac{N_{e}}{10^{-3}\text{cm}^{-3}}\right)\left(\frac{E}{\text{GeV}}\right)^{-1}\text{kpc}^{-1},\\ \Delta_{\text{QED}}&=4.1\times 10^{-9}\left(\frac{E}{\text{GeV}}\right)\left(\frac{B_{\perp}}{\mu\text{G}}\right)^{2}\text{kpc}^{-1},\\ \Delta_{a}&=-7.8\times 10^{-2}\left(\frac{m_{a}}{\text{neV}}\right)^{2}\left(\frac{E}{\text{GeV}}\right)^{-1}\text{kpc}^{-1},\\ \Delta_{\gamma a}&=1.52\times 10^{-2}\left(\frac{g_{a\gamma\gamma}}{10^{-11}\text{GeV}^{-1}}\right)\left(\frac{B_{\perp}}{\mu\text{G}}\right)\text{kpc}^{-1}.\\ \end{split}$ (4)
@@
Now the energy variation in the polarization is suppressed in the region with $E\gg E_{c}$, but if $E\ll E_{c}$ leads to an undetectably small polarization signal overall, we might expect that $E_{c}$ must lie within some specific range of energies (comparable to those we observe) in order for the energy variation to be measurable. In this case, the sensitivity region will be bounded below in $g_{a\gamma\gamma}$ by the requirement that $l_{\text{osc}}\lesssim d$, and to the sides by $g_{a\gamma\gamma}\propto m_{a}^{2}$ lines of fixed $E_{c}$. The sensitivity region could potentially be improved in this case by combining AMS-02 polarization measurements with those from lower-energy instruments, but we will not pursue this idea further in this work.
@@
These arguments also clarify how measurements of gamma-ray polarization could improve on previous studies in the X-ray band; detection of polarization is easiest for $E\gtrsim E_{c}$, with detection of energy dependence requiring $E\sim E_{c}$, so measuring polarization at higher $E$ gives access to parameter space that predicts higher values of $E_{c}$. Since where the plasma mass can be neglected, $E_{c}\propto m_{a}^{2}/g_{a\gamma\gamma}$, this implies we will be able to test higher $m_{a}$ for a given value of $g_{a\gamma\gamma}$; the high-$m_{a}$ cutoff associated with the critical energy will move to the right, proportionally to $\sqrt{E}$. Thus we expect that moving from keV-scale X-ray polarization to GeV-scale gamma-ray polarization could buy roughly three orders of magnitude of reach in $m_{a}$. Thus for example Ref. [11] forecasts that X-ray polarization studies of the bright extragalactic source NGC1275 will have sensitivity down to $g_{a\gamma\gamma}$ slightly below $10^{-12}\text{GeV}^{-1}$, up to masses around $m_{a}\sim 10^{-12}$ eV; we may hope for a comparable sensitivity up to $m_{a}\sim 10^{-9}$ eV, although the sensitivity to $g_{a\gamma\gamma}$ will depend on our ability to measure small polarization fractions and hence on the photon statistics.
@@ II.4 Selecting and modeling Galactic sources
In this work, we will assume supernova remnants are unpolarized sources, and consider pulsars both for the polarized and unpolarized cases. We employ the same sample of six pulsars used in Ref. [15]. Note that in order to obtain a large oscillation signal, it is favorable to have a large $B_{\perp}$ component and a long path length. We observe that the radial component of the total $B$-field is relatively small compared with the azimuthal and perpendicular components. Accordingly, we prefer sources at low Galactic latitudes where the emitted photons can pass through the spiral arms, and sources that have relatively large and well-determined distances. The sample we use corresponds to six bright pulsars selected according to these criteria. We also selected four of the five brightest supernova remnants from the Fermi-LAT fourth source catalog (4FGL) [44] using the same criteria; we did not include the W28 supernova remnant, which is the fourth brightest overall, due to a relatively small transverse magnetic field along the line of sight to this source. The distances to these sources are taken from Refs. [16, 45]. [^2] The uncertainties on the distances to IC443 and W44 are not clearly defined in the original references (although they are certainly non-negligible). The distance to W51C is taken to be 5.5kpc, following Ref. [16], which is also consistent with the measured value of $5.41^{+0.31}_{-0.28}$kpc given in Ref. [46]. The positions of these sources are listed in Table 1 and are shown, together with the magnetic field strength in the Galactic plane, in Fig. 1. The exposure for AMS-02 [24] can be reduced in regions such as the North Pole and the South Atlantic Anomaly due to a decrease in data acquisition efficiency. The sources considered in this paper have been verified to be located outside these regions of low exposure.
@@
Figure 1: The strength of the Galactic magnetic field, based on the model in Ref. [33], as a function of position in the Galactic plane. The selected pulsars and supernova remnants are indicated as crosses and triangles, respectively. The position of the sun is marked as a red point.
The pulsars are fairly young and rotation powered. Their emission spectra are modelled as a power law with an exponential cutoff as in Ref. [15]:
@@
Pulsar name $N_{0}$ $E_{0}$ $\Gamma$ $E_{\text{cut}}$
@@
The emission spectra of the supernova remnants are modelled by a log-normal representation as in the Fourth Fermi-LAT catalog [44],
@@
SNR name $K$ $E_{0}$ $\alpha$ $\beta$
@@ II.5 Selecting and modeling extragalactic sources
NGC1275 has a Galactic longitude of $150.58^{\circ}$ and a Galactic latitude of $-13.26^{\circ}$; its estimated distance is 68.2 Mpc. The emission spectrum of NGC1275 is also modelled by a log-normal representation (as in Eq. 14) with the best fit parameters [44] again given in Table 3. We follow Ref. [11] for the modeling of the magnetic field. Specifically, we take a central magnetic field strength of $25\mu$G as indicated in Ref. [47]. The magnetic field falls off radially as $B\propto n_{e}^{0.7}$ as modeled in Ref. [48] for galaxy clusters. The radial behavior of electron density is modeled in Ref. [49] as
@@
There is not much known on the coherence length of the Perseus cluster, so we parametrize it using the values motivated by the study on the structure of the magnetic field in the cool core cluster A2199 [50]. The minimum coherence length is assumed to be 3.5kpc and the maximum coherence length 10kpc. The magnetic field along the line of sight is simulated with the above setup; a set of coherence lengths is drawn randomly from the probability distribution $P(l=x)\propto x^{-1.2}$, until the sum of the lengths adds up to 1 Mpc. In each domain, the magnetic field is assumed to be constant with fixed random direction.
@@ II.6 Expected perturbation of the gamma-ray polarization
Figure 2: Example of expected photon intensity and polarization from the pulsar PSR J2021+3651, with ALP parameters $g_{a\gamma\gamma}=3.543\times 10^{-10}$ GeV-1, $m_{a}=4.41\times 10^{-9}$ eV. (a) The intensity of photons at 3 GeV versus the distance along the line of sight from the pulsar. (b) The intensity of photons versus the photon energy. (c) The polarization degree of photons at 3 GeV versus the distance along the line of sight from the pulsar. (d) The polarization degree of photons versus the photon energy. (e) The polarization angle of photons at 3 GeV versus the distance along the line of sight from the pulsar. (f) The polarization angle of photons versus the photon energy, at the Earth’s location. The red curve assumes an initially unpolarized source; the blue and violet curves assume initially 100% polarized source with different initial angles (see text for details). The transversal magnetic field $\mathrm{B_{\perp}}$ in units of $\mu$G is shown as the dashed green line in (a), (c), and (e).
@@ III Polarization measurement by pair production
Figure 3: Kinematics of electron-positron pair production. The $x$ and $y$ axes are aligned with the directions of the electric field and magnetic field carried by the light source, respectively. The 3-momenta of the incoming photon, the emitted electron and positron are denoted as $\vec{k}_{\gamma}$, $\vec{p}_{e^{-}}$, and $\vec{p}_{e^{+}}$. The $x-$ or electric field direction is also called the polarization direction of the source. The electron and positron form a plane (green) which is perpendicular to the $x-y$ plane (blue). The azimuthal angle $\phi$ is defined as the angle between the electron-positron plane and the direction of the electric field.
@@
where $A$ is the analyzing power depending on the angular resolution of the detector and the kinematics of the pair production, and $P$ is the degree of the photon polarization. For AMS-02, the value of $A$ in the GeV energy range can be approximated as a constant value 0.14 [53]. We will assume future successor instruments will have a similar value of $A$ and will differ only in their acceptance.
@@
The measurement of gamma rays with the AMS-02 experiment is described in Ref. [24]. The effective acceptance using photons converted in the upper detector increases as the energy increases from nearly zero at $\sim 100$ MeV, reaching 140 cm2sr at $\sim$ 5 GeV, and then decreases to zero at 1 TeV. For photon sources close to the zenith of AMS-02, the effective area is $\sim$ 180 cm2. The proposed next generation magnetic spectrometer AMS-100 will increase the acceptance of gamma rays by at least a factor of $\sim$ 100 [26]. Another similar proposal [27] intends to increase the acceptance by one order of magnitude compared to AMS-02.

### v2

@@ Abstract
We study the detectability of GeV-band gamma-ray polarization with the AMS-02 experiment and its proposed successor AMS-100, from Galactic and extragalactic sources. Characterizing gamma-ray polarization in this energy range could shed light on gamma-ray emission mechanisms in the sources; physics beyond the Standard Model, such as the presence of axion-like particles (ALPs), could also induce a distinctive energy-dependent polarization signal due to propagation effects in magnetic fields. We present estimates for the minimum detectable polarization from bright sources and the forecast reach for axion-like particles (ALPs). We show that AMS-02 will have sensitivity to gamma-ray polarization only for the brightest steady-state Galactic sources, such as the Vela and Geminga pulsars; it is not expected to be capable of detecting polarization in Galactic or extragalactic sources that have been previously proposed as good targets for ALP searches with gamma-ray intensity measurements. However, AMS-100 observing the extragalactic source NGC1275 would be expected to probe new parameter space even for unfavorable B-field models, with prospects to measure the energy-dependence of such a signal. For Galactic sources, polarization measurements could provide a unique test of scenarios where ALPs induce energy-dependent features in the photon intensity. However, in the absence of a bright transient source (such as a Galactic supernova), the parameter space that would be probed by this approach with ten years of AMS-100 data is already nominally excluded by other experiments, although this conflict may be avoided in specific ALP models.
@@ I Introduction
The existence of dark matter has long been one of the greatest puzzles in our understanding of the Universe. One of the leading candidates for dark matter particles is a hypothetical particle called the axion [1, 2, 3, 4, 5] introduced by Peccei and Quinn [6, 7] to resolve the puzzle of the lack of observed CP violation in quantum chromodynamics (QCD). More generally, axion-like particles (ALPs) are ultralight pseudoscalar bosons predicted in many theoretical frameworks beyond the Standard Model of particle physics, e.g., [8, 9, 10]. In contrast to QCD axions, the characteristic parameters of ALPs—the mass $m_{a}$ and the photon-axion coupling $g_{a\gamma\gamma}$—are not necessarily related to each other. In the presence of an external magnetic field, the photon-ALP mixing can lead to energy-dependent modifications to both the flux and the polarization of photons from astrophysical sources, over a wide range of frequencies.
The effects of ALPs on the polarization of light have been studied for various astrophysical sources (e.g., gamma-ray bursts [11], magnetic white dwarfs [12], active galactic nuclei [13, 14], and pulsars [15]). These studies are typically applied to photon energies below the MeV scale. Extending the studies of polarization to the GeV region requires the measurement of electron-positron pair production. To date, no data on polarization has been released for photons of energies above the pair creation threshold ($\sim$ 1 MeV), although the Fermi Large Area Telescope (Fermi-LAT) has the potential to detect high degrees of linear polarization from some of the brightest gamma-ray sources [16]. Challenges and future directions for gamma-ray polarimetry have been reviewed in Ref. [17].
In the gamma-ray band, ALP-induced modification of intensity spectra from Galactic pulsars [18] and supernova remnants [19] has already been studied in data from the Fermi-LAT gamma-ray telescope. Intriguingly, these studies have claimed a hint of a possible signal, with a favored mass of $\sim$ neV and coupling $g_{a\gamma\gamma}\sim 10^{-10}$ GeV-1. The tension of these results with the limits from the CAST helioscope [20] might be reconciled by considering the different environmental conditions inside the Sun and the dilute interstellar medium where the photon-axion mixing occurs [21], or by non-standard ALP couplings [22]. Additional experimental data such as polarization from these Galactic sources could provide valuable input in testing the ALP hypothesis and perhaps probing the properties of ALPs.
@@
In this paper, we estimate the possible imprint of ALP dark matter in energy-dependent gamma-ray polarization, and discuss the implications for the sensitivity of AMS-02 and proposed next generation spectrometer experiments in space. Proposals for such experiments include ALADInO and AMS-100  [29, 30]; we will focus on AMS-100, which aims to have near-$4\pi$ sr sky coverage, and geometric acceptance exceeding that of AMS-02 by roughly a factor of 1000. We consider the extragalactic source NGC1275 at the center of the Perseus cluster, which has been studied in the context of ALP searches using X-ray polarimetry (e.g. in Ref. [14]) and gamma-ray intensity measurements (e.g. [31, 21, 32]), and a range of bright Galactic pulsars and supernovae.
@@ II.1 Photon-ALP mixing
[…]
@@
[…]
@@
Now the energy variation in the polarization is suppressed in the region with $E\gg E_{c}$, but if $E\ll E_{c}$ leads to an undetectably small polarization signal overall, we might expect that $E_{c}$ must lie within some specific range of energies (comparable to those we observe) in order for the energy variation to be measurable. In this case, the sensitivity region will be bounded below in $g_{a\gamma\gamma}$ by the requirement that $l_{\text{osc}}\lesssim d$, and to the sides by $g_{a\gamma\gamma}\propto m_{a}^{2}$ lines of fixed $E_{c}$. The sensitivity region could potentially be improved in this case by combining gamma-ray polarization measurements with those from lower-energy instruments, but we will not pursue this idea further in this work.
@@
These arguments also clarify how measurements of gamma-ray polarization could improve on previous studies in the X-ray band; detection of polarization is easiest for $E\gtrsim E_{c}$, with detection of energy dependence requiring $E\sim E_{c}$, so measuring polarization at higher $E$ gives access to parameter space that predicts higher values of $E_{c}$. Since where the plasma mass can be neglected, $E_{c}\propto m_{a}^{2}/g_{a\gamma\gamma}$, this implies we will be able to test higher $m_{a}$ for a given value of $g_{a\gamma\gamma}$; the high-$m_{a}$ cutoff associated with the critical energy will move to the right, proportionally to $\sqrt{E}$. Thus we expect that moving from keV-scale X-ray polarization to GeV-scale gamma-ray polarization could buy roughly three orders of magnitude of reach in $m_{a}$. Thus for example Ref. [14] forecasts that X-ray polarization studies of the bright extragalactic source NGC1275 will have sensitivity down to $g_{a\gamma\gamma}$ slightly below $10^{-12}\text{GeV}^{-1}$, up to masses around $m_{a}\sim 10^{-12}$ eV (where there is a sharp cutoff in sensitivity); we may hope for a comparable sensitivity up to $m_{a}\sim 10^{-9}$ eV, although the sensitivity to $g_{a\gamma\gamma}$ will depend on our ability to measure small polarization fractions and hence on the photon statistics.
@@ II.4 Selecting and modeling Galactic sources
In this work, we will assume supernova remnants are unpolarized sources, and consider pulsars both for the polarized and unpolarized cases. We employ the same sample of six pulsars used in Ref. [18]. Note that in order to obtain a large oscillation signal, it is favorable to have a large $B_{\perp}$ component and a long path length. We observe that the radial component of the total $B$-field is relatively small compared with the azimuthal and perpendicular components. Accordingly, we prefer sources at low Galactic latitudes where the emitted photons can pass through the spiral arms, and sources that have relatively large and well-determined distances. The sample we use corresponds to six bright pulsars selected according to these criteria. We also selected four of the five brightest supernova remnants from the Fermi-LAT fourth source catalog (4FGL) [47] using the same criteria; we did not include the W28 supernova remnant, which is the fourth brightest overall, due to a relatively small transverse magnetic field along the line of sight to this source. The distances to these sources are taken from Refs. [19, 48]. In addition to these sources, which are chosen to be favorable for observations of ALP-photon oscillation, we also consider two sources chosen purely for their brightness: the Vela and Geminga pulsars. These pulsars are very nearby and so are unlikely to be good candidates for observing ALP-induced polarization, but their high photon fluxes make them good candidates for observations of intrinsic (astrophysical) polarization. [^2] The uncertainties on the distances to IC443 and W44 are not clearly defined in the original references (although they are certainly non-negligible). The distance to W51C is taken to be 5.5kpc, following Ref. [19], which is also consistent with the measured value of $5.41^{+0.31}_{-0.28}$kpc given in Ref. [49].
The positions of these sources are listed in Table 1 and are shown, together with the magnetic field strength in the Galactic plane, in Fig. 1 (positions for Geminga and Vela are not shown because they essentially coincide with the Sun’s location on the scale of this plot). The exposure for AMS-02 [27] can be reduced in regions such as the North Pole and the South Atlantic Anomaly due to a decrease in data acquisition efficiency. The sources considered in this paper have been verified to be located outside these regions of low exposure.
@@
Geminga 195.13 4.27 0.25
Vela 263.55 -2.79 0.29
Figure 1: The strength of the Galactic magnetic field, based on the model in Ref. [36], as a function of position in the Galactic plane. The selected pulsars and supernova remnants are indicated as crosses and triangles, respectively. The position of the Sun is marked as a red point. Geminga and Vela are not indicated because of their proximity to the Sun (which also makes them poor targets for an ALP search).
The pulsars are fairly young and rotation powered. Their emission spectra are modeled as a power law with an exponential cutoff as in Ref. [18]:
@@
Pulsar name $N_{0}$
$10^{-9}\rm{MeV}^{-1}\rm{cm}^{-2}\rm{s}^{-1}]$
$E_{0}$
$\rm{GeV}]$
$\Gamma$ $E_{\text{cut}}$
$\rm{GeV}]$
@@
The emission spectra of the supernova remnants are modeled by a log-normal representation as in the Fourth Fermi-LAT catalog [47],
@@
SNR name $K$
$10^{-9}\rm{MeV}^{-1}\rm{cm}^{-2}\rm{s}^{-1}]$
$E_{0}$
$\rm{GeV}]$
$\alpha$ $\beta$
@@
Unlike the other pulsars, Geminga and Vela have significantly curved spectra and are described by a sub-exponentially cutoff power law,
$\frac{dN}{dE}=N_{0}\left(\frac{E}{E_{0}}\right)^{\gamma_{0}+d/b}\exp\left[\frac{d}{b^{2}}\left(1-\left(\frac{E}{E_{\text{cut}}}\right)^{b}\right)\right].$ (15)
For Geminga, we have a prefactor $N_{0}=0.30696\times 10^{-9}\rm{MeV}^{-1}\rm{cm}^{-2}\rm{s}^{-1}$, local spectral index $\gamma_{0}=-2.0288$, local curvature $d=0.71448$, index $b=0.6806$, and energy scale $E_{0}=1.70533$ GeV. For Vela, we have $N_{0}=0.41083\times 10^{-9}\rm{MeV}^{-1}\rm{cm}^{-2}\rm{s}^{-1}$, local spectral index $\gamma_{0}=-2.2607$, local curvature $d=0.57942$, index $b=0.4922$, and energy scale $E_{0}=1.97637$ GeV.
@@ II.5 Selecting and modeling extragalactic sources
[…]
@@
There is not much known on the coherence length of the B-field of the Perseus cluster, so we parametrize it (as in Ref. [14]) using the values motivated by a study of the structure of the magnetic field in the cool core cluster A2199 [53]. The minimum coherence length is assumed to be 3.5kpc and the maximum coherence length 10kpc. The magnetic field along the line of sight is simulated with the above setup; a set of coherence lengths is drawn randomly from the probability distribution $P(l=x)\propto x^{-1.2}$, until the sum of the lengths adds up to 1 Mpc. In each domain, the magnetic field is assumed to be constant with fixed random direction.
@@ II.6 Expected perturbation of the gamma-ray polarization
[…]
@@ III Polarization measurement by pair production
Figure 3: Kinematics of electron-positron pair production. The $x$ and $y$ axes are aligned with the directions of the electric field and magnetic field carried by the light source, respectively. The 3-momenta of the incoming photon, the emitted electron and positron are denoted as $\vec{k}_{\gamma}$, $\vec{p}_{e^{-}}$, and $\vec{p}_{e^{+}}$. The $x$- or electric field direction is also called the polarization direction of the source. The electron and positron form a plane (green) which is perpendicular to the xyx-y plane (blue). The azimuthal angle $\phi$ is defined as the angle between the electron-positron plane and the direction of the electric field.
@@
where $A$ is the analyzing power depending on the angular resolution of the detector and the kinematics of the pair production, and $P$ is the degree of the photon polarization. For silicon-strip pair-production detectors operating in the energy range above 0.1 GeV, such as AMS-02, the value of $A$ can be approximated as a constant value 0.14 [57, 28]. We will assume future successor instruments will have a similar value of $A$ and will differ only in their acceptance. That said, we caution that the effective polarization asymmetry can depend sensitively on the design of the detector; Ref. [28] finds that the effective value of $A$ is significantly suppressed for the Fermi-LAT, for example. While that work argues that the AMS-02 tracker has a number of properties that are favorable for gamma-ray polarimetry (specifically, “thin wafers, narrow readout pitch, and large distance between layers”), these properties would not necessarily be preserved in AMS-100. We explore the potential of AMS-100 for gamma-ray polarimetry with this assumption ($A\simeq 0.14$) in part to understand the science case for taking polarimetry into account in the design of AMS-100.
@@
The measurement of gamma rays with the AMS-02 experiment is described in Ref. [27]. The effective acceptance using photons converted in the upper detector increases as the energy increases from nearly zero at $\sim 100$ MeV, reaching 140 cm2sr at $\sim$ 5 GeV, and then decreases to zero at 1 TeV. For photon sources close to the zenith of AMS-02, the effective area is $\sim$ 180 cm2. For point sources, it is the exposure rather than the overall acceptance that is important. A realistic calculation of the AMS-02 exposure over six years, accounting for the efficiency of various selection cuts, is conducted in Ref. [27] for one specific energy (2 GeV). For our AMS-02 calculations, we digitize the results of that work for the exposure (for the conversion analysis, not the calorimeter events) at the positions of our sources. We then rescale the results by a factor of $10/6$ to obtain a 10-year sensitivity. The resulting exposures (at 2 GeV) are described in Table 4. We extend the results to our full energy range (100 MeV – 10 GeV) by assuming the energy dependence of the exposure is dominated by the energy dependence of the effective area, which is given for the conversion analysis in that same work.
Table 4: Estimated signal and background (bkg) counts for AMS-02 in ten years for selected sources, along with exposure at 2 GeV. The first six entries correspond to pulsars, the next four to supernovae, and the last three to our example extragalactic source and Geminga and Vela pulsars.
Source name Estimated counts for AMS-02 in ten years Estimated bkg counts for AMS-02 in ten years Exposure at 2 GeV  $10^{9}$ cm2 s]
J1420-6048 14 62 0.308
J1648-4611 6 194 0.729
J1702-4128 8 250 0.866
J1718-3825 29 200 0.923
J2021+3651 246 173 1.79
J2240+5832 13 71 1.70
IC443 185 49 1.60
W44 209 357 1.43
W51C 124 219 1.52
W49B 53 272 1.48
NGC1275 133 13 1.82
Geminga 2244 13 1.62
Vela 2191 12 0.737
The proposed next-generation magnetic spectrometer AMS-100 aims for an effective acceptance of 30 m2sr for photon conversions in the silicon tracker layers [29]. Compared to the peak effective acceptance for AMS-02 conversion events, which is 140 cm2sr as discussed above, we expect an increase in statistics by roughly a factor of 2000, averaged over the whole sky.
[…]
Ref. [29] also suggests a smaller AMS-100 pathfinder. The pathfinder would have geometrical acceptance of $0.2\times$ the AMS-02 value, so we simply scale down the exposure by this factor. (Note that Ref. [29] characterizes this pathfinder as $10\%$ scale”, but roughly a factor of 2 of this loss would come from removing the calorimeter, which is irrelevant for our analysis.) We hereafter denote this $20\%$-scale pathfinder as AMS-100P.
Having obtained the energy-dependent exposure for both telescopes, we integrate these functions over the energy spectra of the sources (as given in Tables 2-3) to obtain the total number of expected counts.